A certain journey took 40minutes to complete. If half the distance was travelled at an average speed of 100 km/h and the other half was at an average speed of 60 km/h what was the length of the journey.
Answer by lwsshak3(11628)

A certain journey took 40minutes to complete. If half the distance was travelled at an average speed of 100 km/h and the other half was at an average speed of 60 km/h what was the length of the journey.
***
let x=length of the journey
travel time=distance/speed
100/(x/2)+60/(x/2)=40/60
200/x+120/x=2/3
lcd:3x
600+360=2x
960=2x
x=480
length of the journey=480 km